# X2-Gaussian: 4D Radiative Gaussian Splatting for Continuous-time Tomographic Reconstruction

### v1

@@ Abstract
Four-dimensional computed tomography (4D CT) reconstruction is crucial for capturing dynamic anatomical changes but faces inherent limitations from conventional phase-binning workflows. Current methods discretize temporal resolution into fixed phases with respiratory gating devices, introducing motion misalignment and restricting clinical practicality. In this paper, We propose X2-Gaussian, a novel framework that enables continuous-time 4D-CT reconstruction by integrating dynamic radiative Gaussian splatting with self-supervised respiratory motion learning. Our approach models anatomical dynamics through a spatiotemporal encoder-decoder architecture that predicts time-varying Gaussian deformations, eliminating phase discretization. To remove dependency on external gating devices, we introduce a physiology-driven periodic consistency loss that learns patient-specific breathing cycles directly from projections via differentiable optimization. Extensive experiments demonstrate state-of-the-art performance, achieving a 9.93 dB PSNR gain over traditional methods and 2.25 dB improvement against prior Gaussian splatting techniques. By unifying continuous motion modeling with hardware-free period learning, X2-Gaussian advances high-fidelity 4D CT reconstruction for dynamic clinical imaging. Project website at: https://x2-gaussian.github.io/.
@@ 1 Introduction
Four-dimensional computed tomography (4D CT) has become a cornerstone in dynamic medical imaging, especially for respiratory motion management in clinical applications such as image-guided radiotherapy (IGRT) [12, 38]. By capturing both spatial and temporal information of the chest cavity during breathing cycles, 4D CT enables clinicians to monitor and assess respiratory-induced tumor motion and other dynamic anatomical changes during treatment [49, 13, 3].
@@ 2.1 CT Reconstruction
Traditional 3D computed tomography reconstruction methods mainly include two categories: analytical algorithms [15, 55] and iterative algorithms [1, 46, 35, 44]. Analytical methods estimate the radiodensity by solving Radon transformation and its inverse version. Iterative algorithms are based on optimization over iterations. In recent years, deep learning based models [2, 18, 26, 32, 34] like CNNs have been employed to learn a brute-force mapping from X-ray projections to CT slices. Another technical route is to employ the 3D rendering algorithms such as neural radiance fields (NeRF) [37] and 3D Gaussian Splatting (3DGS) [29] to solve the CT reconstruction prolblem in a self-supervised manner, i.e. using only 2D X-rays for training. Based on these algorithms, when coping with 4D CTs, researchers typically segment the projections into ten discrete respiratory phases for sequential 3D reconstruction. This approach not only necessitates external devices for phase measurement during scanning but also impedes accurate modeling of the continuous motion of anatomical structures. Concurrent work [17] also employs dynamic Gaussian splatting. However, they merely establish ten timestamps corresponding to ten phases, thereby maintaining a discrete representation. In contrast, this paper is dedicated to achieving truly continuous-time 4D CT reconstruction.
@@ 4.3 Self-Supervised Respiratory Motion Learning
  • Log-Space Parameterization: We represent $\hat{T}=\exp(\hat{\tau})$ where $\hat{\tau}\in\mathbb{R}$ is an unbounded learnable variable. This ensures positivity and provides smoother gradient updates compared to direct period estimation. This logarithmic parameterization ensures $T$ remains positive, improves numerical stability by preventing extremely small period values, and creates a more uniform gradient landscape for optimization.
@@ 7 Details of Dataset
We collected 4D CT scans from the DIR dataset [9], which were acquired from patients with malignant thoracic tumors (esophageal or lung cancer). Each 4D CT was divided into 10 3D CT volumes based on respiratory signals captured by a real-time position management respiratory gating system [28]. For each patient, the CT dimensions are $256\times 256$ in the x and y axes, while the z-axis dimension varies from $94$ to $112$ slices. The z-axis resolution is $2.5$ $mm$, and the xy-plane resolution ranges between $0.97$ and $1.16$ $mm$. The CT scan coverage encompasses the entire thoracic region and upper abdomen. Following the approach in literature [59, 7], we preprocessed the original data by normalizing the density values to the range of $[0,1]$. We simulated the classical one-minute sampling protocol used in clinical settings by uniformly sampling 300 paired time points and angles within a one-minute duration and a 00 to $360$ angular range. Based on the respiratory phase corresponding to each timestamp, we selected the appropriate 3D CT volume, and then utilized the tomographic imaging toolbox TIGRE [4] to capture $512\times 512$ projections.
@@
4D CTs in 4DLung dataset [23] were collected from non-small cell lung cancer patients during their chemoradiotherapy treatment. All scans were respiratory-synchronized into 10 breathing phases. For each patient, the CT scans have dimensions of $512\times 512$ pixels in the transverse plane, with the number of axial slices varying between $91$ and $135$. The spatial resolution is $0.9766$ to $1.053$ $mm$ in the transverse plane and $3$ $mm$ in the axial direction. Following the same pipeline as DIR dataset, We captured 300 projections with sizes of $1024\times 1024$.
@@
The 4D CT images from the SPARE dataset [45] have dimensions of $450\times 450$ pixels in the transverse plane and $220$ slices in the axial direction, with an isotropic spatial resolution of $1.0$ $mm$ in all directions. Following the same methodology as the DIR dataset, we acquired 300 projections, each with dimensions of $512\times 512$ pixels.

### v2

@@ Abstract
Four-dimensional computed tomography (4D CT) reconstruction is crucial for capturing dynamic anatomical changes but faces inherent limitations from conventional phase-binning workflows. Current methods discretize temporal resolution into fixed phases with respiratory gating devices, introducing motion misalignment and restricting clinical practicality. In this paper, We propose X2-Gaussian, a novel framework that enables continuous-time 4D-CT reconstruction by integrating dynamic radiative Gaussian splatting with self-supervised respiratory motion learning. Our approach models anatomical dynamics through a spatiotemporal encoder-decoder architecture that predicts time-varying Gaussian deformations, eliminating phase discretization. To remove dependency on external gating devices, we introduce a physiology-driven periodic consistency loss that learns patient-specific breathing cycles directly from projections via differentiable optimization. Extensive experiments demonstrate state-of-the-art performance, achieving a 9.93 dB PSNR gain over traditional methods and 2.25 dB improvement against prior Gaussian splatting techniques. By unifying continuous motion modeling with hardware-free period learning, X2-Gaussian advances high-fidelity 4D CT reconstruction for dynamic clinical imaging. Code is publicly available at: https://x2-gaussian.github.io/.
@@ 1 Introduction
Four-dimensional computed tomography (4D CT) has become a cornerstone in dynamic medical imaging[24, 55, 56, 25], especially for respiratory motion management in clinical applications such as image-guided radiotherapy (IGRT) [12, 38]. By capturing both spatial and temporal information of the chest cavity during breathing cycles, 4D CT enables clinicians to monitor and assess respiratory-induced tumor motion and other dynamic anatomical changes during treatment [49, 13, 3].
@@ 2.1 CT Reconstruction
Traditional 3D computed tomography reconstruction methods mainly include two categories: analytical algorithms [15, 57] and iterative algorithms [1, 46, 35, 44]. Analytical methods estimate the radiodensity by solving Radon transformation and its inverse version. Iterative algorithms are based on optimization over iterations. In recent years, deep learning based models [2, 18, 26, 63, 32, 34] like CNNs have been employed to learn a brute-force mapping from X-ray projections to CT slices. With the development of 3D deep learning techniques[37, 29, 60, 58, 61, 59, 62, 36, 53, 65], another technical route is to employ the 3D rendering algorithms such as neural radiance fields (NeRF) [37] and 3D Gaussian Splatting (3DGS) [29] to solve the CT reconstruction prolblem in a self-supervised manner, i.e. using only 2D X-rays for training. Based on these algorithms, when coping with 4D CTs, researchers typically segment the projections into ten discrete respiratory phases for sequential 3D reconstruction. This approach not only necessitates external devices for phase measurement during scanning but also impedes accurate modeling of the continuous motion of anatomical structures. Concurrent work [17] also employs dynamic Gaussian splatting. However, they merely establish ten timestamps corresponding to ten phases, thereby maintaining a discrete representation. In contrast, this paper is dedicated to achieving truly continuous-time 4D CT reconstruction.
@@ 4.3 Self-Supervised Respiratory Motion Learning
[…]
@@ 7 Details of Dataset
We collected 4D CT scans from the DIR dataset [9], which were acquired from patients with malignant thoracic tumors (esophageal or lung cancer). Each 4D CT was divided into 10 3D CT volumes based on respiratory signals captured by a real-time position management respiratory gating system [28]. For each patient, the CT dimensions are $256\times 256$ in the x and y axes, while the z-axis dimension varies from $94$ to $112$ slices. The z-axis resolution is $2.5$ mmmm, and the xy-plane resolution ranges between $0.97$ and $1.16$ mmmm. The CT scan coverage encompasses the entire thoracic region and upper abdomen. Following the approach in literature [67, 7], we preprocessed the original data by normalizing the density values to the range of $[0,1]$. We simulated the classical one-minute sampling protocol used in clinical settings by uniformly sampling 300 paired time points and angles within a one-minute duration and a 0 to $360$ angular range. Based on the respiratory phase corresponding to each timestamp, we selected the appropriate 3D CT volume, and then utilized the tomographic imaging toolbox TIGRE [4] to capture $512\times 512$ projections.
@@
4D CTs in 4DLung dataset [23] were collected from non-small cell lung cancer patients during their chemoradiotherapy treatment. All scans were respiratory-synchronized into 10 breathing phases. For each patient, the CT scans have dimensions of $512\times 512$ pixels in the transverse plane, with the number of axial slices varying between $91$ and $135$. The spatial resolution is $0.9766$ to $1.053$ mmmm in the transverse plane and 33 mmmm in the axial direction. Following the same pipeline as DIR dataset, We captured 300 projections with sizes of $1024\times 1024$.
@@
The 4D CT images from the SPARE dataset [45] have dimensions of $450\times 450$ pixels in the transverse plane and $220$ slices in the axial direction, with an isotropic spatial resolution of $1.0$ mmmm in all directions. Following the same methodology as the DIR dataset, we acquired 300 projections, each with dimensions of $512\times 512$ pixels.